# DSSYK at Infinite Temperature: The Flat-Space Limit and the ‘t Hooft Model

### v1

@@
DSSYK at Infinite Temperature: The Flat-Space Limit and the ‘t Hooft Model
@@ 1 The Holographic Principle
A naive version of the Holographic Principle states that a region of space $\Gamma$ and everything in it can be described by a quantum system with a Hilbert space of dimension $\exp{A/4G},$ $A$ being the area of the boundary of $\Gamma$; and a Hamiltonian acting on that space [1][2]. A more precise version [3] goes as follows:
@@ 1.1 The Flat-Space Limit
The flat-space limit of AdS/CFT is well known [8][9]. It is the limit $N\to\infty$ with the CFT gauge coupling $g_{ym}$ held fixed [^1] This limit is not the standard ‘t Hooft limit, $g_{ym}^{2}N$ held fixed., and energies scaling like $N^{1/4}$ in units of the AdS radius. The corresponding flat-space limit of de Sitter space is similar. It is the limit of large entropy in which the de Sitter radius of curvature goes to infinity and spacetime becomes locally flat. The static patch, out to arbitrary distances from the horizon, becomes flat Rindler space.
@@ 2 DSSYK
The sum in (2.2) is over all $1\leq i_{1}<i_{2}...\leq N$ and $1\leq j_{1}<j_{2}...\leq N.$ The coupling constants $J^{j_{1}j_{2}\cdots j_{p}}_{i_{1}i_{2}\cdots i_{p}},$ which will collectively be donoted $\{J\},$ are independently drawn from a Gaussian random ensemble with variance
@@ 3 Expectation Values and Ensemble Averages
The ensemble average of a function $F(J)$ is defined as an integral over $J$ weighted by the gaussian measure with variance (2.3). We will denote it by large round brackets, i.e.,
[…]
@@ 3.1 Self-Averaging
for some $\alpha>0.$ In that case ${\cal{O}}$ is self-averaging to order $\alpha.$ One may even find $\text{Var}\langle{\cal{O}}\rangle$ tends to zero faster than any power of $N.$ Then ${\cal{O}}$ is “perfectly self-averaging.” We will see later in the next section that in the double-scaled limit many important quantities are perfectly self averaging.
@@ 4.4 Diagramatic Notation
Figure 1: Fermion and $JJ$ propagators.
@@
The expectation value of $W$ invovles contracting the fermion lines in all possible ways, one of which is shown in figure 3. The red dashed lines are uncontracted and the result is still a function of $\{J\}$.
Figure 3: The expectation value of $W$ is computed by contracting the fermions but leaving the $JJ$ insertions uncontracted.
@@
Figure 4: The final step in computing the average of the expectation value of $W$ is to contract the $JJ$ insertions.
@@
Figure 6: The average of the expectation value of $WW.$
@@
In the double-scaled limit, $N\to\infty$ with $\lambda=p^{2}/N$ fixed, this goes to zero [^2] There are $p$ fermion loops for one $J$ in each copy of $W$ (except for the $J$’s which are contracted within the same copy). This gives a factor $N^{p}/p!$. The fermion loops in the other copy of $W$ does not give another factor, since the $JJ$ contractions between the two copies require both copies to have the same set of fermion indices. So the factor coming from the fermion loops is not large enough to compensate for the small factor (4.20). faster than any power of $N.$
@@
5 ‘t Hooft Model in The Bulk
@@
These considerations lead to the conclusion that the bulk dual of ${\rm DSSYK_{\infty}}$ in the flat-space limit is a $(1+1)$ dimensional $SU(N)\times U(1)$ gauge theory, with a single multiplet of fermions in the fundamental representation. Moreover the flat-space limit requires that we take the $N\to\infty$ limit. Simply put, the bulk dual is the one-flavor ‘t Hooft model [^4] The $N$ species of SYK fermions are often referred to as flavors but it is more correct to think of them as colors. [14]. For brevity we’ll refer to this as $QCD_{2}.$
@@
5.1 The ‘t Hooft and Flat Space Limits
The standard large $N$ limit of gauge theories—say in four dimensions—is the ’t Hooft limit in which the dimensionless gauge coupling $g_{ym}$ goes to zero, with the dimensionless ‘t Hooft coupling [^5] The ‘t Hooft coupling $g_{ym}^{2}N$ is usually denoted by $\lambda.$ To avoid confusion with the ${\rm DSSYK_{\infty}}$ parameter $\lambda$ we denote the ‘t Hooft coupling by $\alpha.$ $\alpha$ remaining finite.
[…]
@@ 5.1 The ‘t Hooft and Flat Space Limits
Something similar is true for $QCD_{2}$ but, because $g_{ym}$ and $\alpha$ are not dimensionless, extra care is needed in defining the limits. We will need to construct dimensionless versions them in order to define the limits correctly.
In $QCD_{2}$ there are two DIMENSIONFUL parameters, the quark mass [^6] We use the notation $m_{q}$ for the shifted (renormalized) quark mass defined in terms of the Lagrangian mass $m_{0}$ by $m_{q}^{2}=m_{0}^{2}-g_{ym}^{2}N/\pi.$ $m_{q}$ and the gauge coupling $g_{ym}.$ Both have units of mass. Naively we might expect the dimensionless coupling parameter to be the ratio $g_{ym}^{2}/m_{q}^{2},$ but in fact things are a bit more subtle  [^7] We are indebted to Ofer Aharony for clarifying discussions and for pointing out an error in the original version of this paper.. To understand why, consider the interaction strength between a quark and antiquark in the singlet state. It is given by single gluon exchange. The $\bar{q}q$ potential is given by,
$V=g_{ym}^{2}N|r|$ (5.2)
where $|r|$ is the distance between $\bar{q}$ and $q$. The quantity
$\tau=g_{ym}^{2}N$ (5.3)
@@
The ’t Hooft Hamiltonian in the lightcone frame is the sum of a free quark term,
$H_{0}=\frac{m_{q}^{2}}{x(1-x)}\ \ \ \ (0<x<1)$ (5.4)
where $x$ is the “longitudinal fraction” of momentum, and a potential term given by (5.2)
$H=H_{0}+V.$ (5.5)
The squared quark mass in the ’t Hooft model must be chosen proportional to $N,$
$m_{q}^{2}=\mu^{2}N\ \ (\mu^{2}\ \text{fixed}).$ (5.6)
@@
$H=m_{q}^{2}\left(\frac{1}{x(1-x)}+\frac{g_{ym}^{2}}{\mu^{2}}|r|\right).$ (5.7)
The dimenionless coupling that remains finite in the $N\to\infty$ limit is the ratio [14],
$\frac{\tau}{m_{q}^{2}}=\frac{g_{ym}^{2}}{\mu^{2}}\equiv\bar{g}^{2}.$ (5.8)
The bar notation in (5.8) is used to represent the dimensionless version of a quantity obtained by dividing by a appropriate power of $\mu.$ $\bar{g}$ is also the single dimensionless parameter of $QCD_{2}$ which is held fixed in the large $N$ flat space limit.
Similarly the ‘t Hooft coupling $\alpha=g_{ym}^{2}N$ is not dimensionless; its dimensionless version is [^8] It is not obvious that (5.9) is the parameter that enters the expansion in (5.10) below. We leave the identification of the expansion parameter for future work. But we expect at least that the expansion is in $(g_{ym}/\bar{\mu})^{2}N$ with some mass-dependent parameter $\bar{\mu}^{2}$ that is kept finite when we take $N\to\infty$. One possibility is that $\bar{\mu}^{2}$ is the shifted mass $m_{q}^{2}=m_{0}^{2}-g_{ym}^{2}N/\pi$ with a requirement that $m_{q}^{2}$ is finite in the large $N$ limit, which implies that the bare quark mass $m_{0}^{2}$ which is of order $N$.,
$\bar{\alpha}=\frac{\alpha}{\mu^{2}}=\bar{g}^{2}N.$ (5.9)
We may now define the flat-space limit and an ‘t Hooft-like limit in terms of $\bar{g}$ and $\bar{\alpha}$ as follows.
@@
‘t Hooft limit:      $N\to\infty,\ \ (\bar{\alpha}\ \ \text{fixed}).$
@@
${\cal{A}}=\sum_{n}\frac{F_{n}(\bar{\alpha})}{N^{n}}$ (5.10)
where each $F_{n}(\bar{\alpha})$ is the sum of diagrams with genus [^9] In pure gauge theory the genus of any diagram is an integer. Quarks introduce boundaries, for example a planar diagram with one quark loop has the topology of a disk which has genus $1/2$ $n/2.$
To take the flat-space limit we first eliminate $N$ by using
$N=\frac{\bar{\alpha}}{\bar{g}^{2}}.$
This gives,
${\cal{A}}=\sum_{n}\frac{F_{n}(\bar{\alpha})}{\bar{\alpha}^{n}}\bar{g}^{2n}.$ (5.11)
The flat-space limit is obtained by letting $\bar{\alpha}\to\infty.$ For this limit to make sense—we assume that it does— the functions
$\frac{F_{n}(\bar{\alpha})}{\bar{\alpha}^{n}}$
must have finite limits [15],
$\lim_{\bar{\alpha}\to\infty}\frac{F_{n}(\bar{\alpha})}{\bar{\alpha}^{n}}={\cal{F}}_{n}.$ (5.12)
@@
${\cal{A}}=\sum_{n}{\cal{F}}_{n}\bar{g}^{2n}.$ (5.13)
Equation (5.13) has the following interpretation: The sum is over worldsheets of genus $n/2.$ The factors ${\cal{F}}_{n}$ are the worldsheet amplitudes for genus $n/2.$ And most interestingly the factor $\bar{g}^{2n}$ is the dependence on the open string coupling constant for a worldsheet of genus $n/2.$
We note that the only dimensionless parameter in $QCD_{2}$ is the open string coupling $\bar{g}.$
@@ 5.2 DSSYK-QCD2 Correspondence
${\rm DSSYK_{\infty}}$ also has a $1/N$ expansion with essentially the same form as (5.10) but with $\bar{\alpha}$ being replaced by $p^{2},$
${\cal{A}}=\sum_{n}\frac{F_{n}(p^{2})}{N^{n}}.$ (5.14)
The duality between ${\rm DSSYK_{\infty}}$ and $QCD_{2}$ requires that we make the identification,
$p^{2}=\bar{\alpha}$
or dividing by $N$ and recalling that $\frac{p^{2}}{N}=\lambda,$ we find the correspondence,
$\lambda=\bar{g}^{2}.$ (5.15)
Thus $\lambda$ is the square of the open string coupling.
@@ 5.3 The Meson Tower
The ’t Hooft model is the large $N$ limit of a gauge theory, but it may also be thought of as an open string theory. The spectrum of mesons in the model consists of a single tower or Regge-like trajectory. Figure 7 is taken from [14] and shows the Regge trajectory for a particular non-zero value of $\mu.$ The horizontal axis is the meson mass squared; the vertical axis is an integer $n\geq 2$ analogous to angular momentum in higher dimensions. Asymptotically for large $n$ the trajectory approaches a straight line.
@@
The operators in (5.16) are in a sense complete in that any gauge invariant operator is a function of them. In addition to the single meson states there are multiple meson states and these span the space of states of the ‘t Hooft model.
@@
6 QCD in Rindler Space and DDSYK

### v3

# DSSYK at Infinite Temperature: The Flat-Space Limit and the ’t Hooft Model

@@
DSSYK at Infinite Temperature: The Flat-Space Limit and the ’t Hooft Model
@@ 1 The Holographic Principle
A naive version of the Holographic Principle states that a region of space $\Gamma$ and everything in it can be described by a quantum system with a Hilbert space of dimension $\exp{A/4G},$ AA being the area of the boundary of $\Gamma$; and a Hamiltonian acting on that space [1][2]. A more precise version [3] goes as follows:
@@ 1.1 The Flat-Space Limit
The flat-space limit of AdS/CFT is well known [8][9]. It is the limit $N\to\infty$ with the CFT gauge coupling $g_{ym}$ held fixed [^1] This limit is not the standard ’t Hooft limit, $g_{ym}^{2}N$ held fixed., and energies scaling like $N^{1/4}$ in units of the AdS radius. The corresponding flat-space limit of de Sitter space is similar. It is the limit of large entropy in which the de Sitter radius of curvature goes to infinity and spacetime becomes locally flat. The static patch, out to arbitrary distances from the horizon, becomes flat Rindler space.
@@ 2 DSSYK
The sum in (2.2) is over all $1\leq i_{1}<i_{2}\cdots\leq N$ and $1\leq j_{1}<j_{2}\cdots\leq N.$ The coupling constants $J^{j_{1}j_{2}\cdots j_{p}}_{i_{1}i_{2}\cdots i_{p}},$ which will collectively be donoted $\{J\},$ are independently drawn from a Gaussian random ensemble with variance
@@ 3 Expectation Values and Ensemble Averages
The ensemble average of a function $F(J)$ is defined as an integral over $J$ weighted by the gaussian measure with variance (2.3). We will denote it by large round brackets [^2] In the previous paper [15] by two of the present authors, ensemble average over $J$ was denoted by $\langle\ \rangle_{J}$., i.e.,
$\Big(F\Big)$ $=$ $\int P(J)F(J)dJ$ (3.2)
@@ 3.1 Self-Averaging
for some $\alpha>0.$ In that case $\cal{O}$} is self-averaging to order $\alpha.$ One may even find $\text{Var}\langle{\cal{O}}\rangle$ tends to zero faster than any power of $N.$ Then $\cal{O}$} is “perfectly self-averaging.” We will see later in the next section that in the double-scaled limit many important quantities are perfectly self-averaging.
@@ 4.4 Diagramatic Notation
Figure 1: Fermion and JJJJ propagators.
@@
The expectation value of $W$ involves contracting the fermion lines in all possible ways, one of which is shown in figure 3. The red dashed lines are uncontracted and the result is still a function of $\{J\}$.
Figure 3: The expectation value of $W$ is computed by contracting the fermions but leaving the JJJJ insertions uncontracted.
@@
Figure 4: The final step in computing the average of the expectation value of $W$ is to contract the JJJJ insertions.
@@
Figure 6: The average of the expectation value of $W$.WW.
@@
In the double-scaled limit, $N\to\infty$ with $\lambda=p^{2}/N$ fixed, this goes to zero [^3] The factors from the summation of fermion indices are not large enough to compensate for the small factor (4.20): There are $p$ fermion loops for one $J$ in each copy of $W$, which gives a factor $N^{p}/p!$; there are also $p$ fermion loops in the other copy of $W$, but they do not give another such factor, since the $JJ$ contraction between the two copies require both copies to have the same set of fermion indices. faster than any power of $N.$
@@
5 ’t Hooft Model in The Bulk
@@
These considerations lead to the conclusion that the bulk dual of $\rm DSSYK_{\infty}$} in the flat-space limit is a $(1+1)$ dimensional $SU(N)\times U(1)$ gauge theory, with a single multiplet of fermions in the fundamental representation. Moreover the flat-space limit requires that we take the $N\to\infty$ limit. Simply put, the bulk dual is the one-flavor ’t Hooft model [^5] The $N$ species of SYK fermions are often referred to as flavors but it is more correct to think of them as colors. [14]. For brevity we’ll refer to this as $QCD_{2}.$
@@
5.1 The ’t Hooft and Flat Space Limits
The standard large $N$ limit of gauge theories—say in four dimensions—is the ’t Hooft limit in which the dimensionless gauge coupling $g_{ym}$ goes to zero, with the dimensionless ’t Hooft coupling [^6] The ’t Hooft coupling $g_{ym}^{2}N$ is usually denoted by $\lambda.$ To avoid confusion with the $\rm DSSYK_{\infty}$} parameter $\lambda$ we denote the ’t Hooft coupling by $\alpha.$ $\alpha$ remaining finite.
$\alpha\equiv g_{ym}^{2}N\ (\alpha\ \text{finite in the 't~Hooft limit})$ (5.1)
@@ 5.1 The ’t Hooft and Flat Space Limits
Something similar is true for $QCD_{2}$ but, because $g_{ym}$ and $\alpha$ are not dimensionless, extra care is needed in defining the limits [^7] We thank Ofer Aharony for helpful comments on this point.. We will need to construct dimensionless versions of them in order to define the limits correctly.
In $QCD_{2}$ there are two DIMENSIONFUL parameters, the quark mass [^8] We use the notation $m_{q}$ for the shifted (renormalized) quark mass defined in terms of the Lagrangian mass $m_{0}$ by $m_{q}^{2}=m_{0}^{2}-g_{ym}^{2}N/\pi.$ $m_{q}$ and the gauge coupling $g_{ym}.$ Both have units of mass. We define the dimensionless coupling parameter $\bar{g}^{2}$ to be the ratio  [^9] This definition of $\bar{g}^{2}$ is different from the one in the previous versions of this paper. Previously, we have defined a barred (dimensionless) quantity by dividing the original quantity by $\mu=m_{q}^{2}/N$, while in the present version we define it by dividing by $m_{q}^{2}$.
$\bar{g}^{2}=\frac{g_{ym}^{2}}{m_{q}^{2}}~.$ (5.2)
The bar notation in (5.2) is used to represent the dimensionless version of a quantity obtained by dividing by an appropriate power of $m_{q}.$
Similarly the ’t Hooft coupling $\alpha=g_{ym}^{2}N$ is not dimensionless; its dimensionless version is [^10] Again, this definition of $\bar{\alpha}$ differs from that in the previous versions of this paper.,
$\bar{\alpha}=\frac{\alpha}{m_{q}^{2}}=\bar{g}^{2}N.$ (5.3)
The ’t Hooft Hamiltonian, which describes mesons in the light cone frame, is the sum of the free quark term and potential term,
$H=H_{0}+V.$ (5.4)
The free quark term is
$H_{0}=\frac{m_{q}^{2}}{x(1-x)}\ \ \ \ (0<x<1),$ (5.5)
where $x$ is the “longitudinal fraction” of momentum.
The potential term results from a single gluon exchange between a quark $q$ and antiquark $\bar{q}$, and is given by
$V=g_{ym}^{2}N|r|,$ (5.6)
where $|r|$ is the distance between $q$ and $\bar{q}$. The quantity
$\tau\equiv g_{ym}^{2}N=\alpha$ (5.7)
@@
$H=m_{q}^{2}\left(\frac{1}{x(1-x)}+{\bar{\alpha}}|r|\right).$ (5.8)
The parameter $\bar{\alpha}$ as defined in (5.3) is the dimenionless coupling which characterizes the interaction. It is usually kept finite in the $N\to\infty$ limit [14].
We may now define the flat-space limit and the ’t Hooft limit in terms of $\bar{g}$ and $\bar{\alpha}$ as follows.
@@
’t Hooft limit: $N\to\infty,\ \ (\bar{\alpha}\ \ \text{fixed}).$
@@
${\cal{A}}=\sum_{n=0}^{\infty}\frac{\tilde{\mathcal{F}}^{(n)}(\bar{\alpha},L)}{N^{n}}$ (5.9)
where each $\tilde{\mathcal{F}}^{(n)}(\bar{\alpha},L)$ is given by the sum of diagrams with genus [^11] In pure gauge theory the genus of any diagram is an integer. Quarks introduce boundaries, for example a planar diagram with one quark loop has the topology of a disk which has genus $1/2.$ $n/2$. Each diagram is IR divergent as in the real QCD, so we need an IR cutoff length $L$ in the expansion (5.9). We define $L$ to be dimensionless, by measuring it relative to $1/\sqrt{\alpha}$.
The functions $\tilde{\mathcal{F}}^{(n)}(\bar{\alpha},L)$ take the form,
$\tilde{\mathcal{F}}^{(n)}(\bar{\alpha},L)$ $=$ $\bar{\alpha}^{n}\times\mathcal{F}^{(n)}(\bar{\alpha},L)$ (5.10)
with each $\mathcal{F}^{(n)}(\bar{\alpha},L)$ being a power series in $\bar{\alpha}$ starting from the zeroth power of $\bar{\alpha}$. (It is a power series in $\bar{\alpha}L^{2}$ for large $L$.) The form (5.10) reflects the fact that an order-$n$ diagram has at least $2n$ vertices (i.e., $n$ factors of $\alpha$[15], and that the power of the cutoff $L$ for the leading diagram for each $n$ does not depend on $n$. An example of this is to increase $n$ by one from a given diagram, by inserting a quark loop into a free gluon propagator: The power of $\alpha$ increases by one, but the infrared behavior of the corrected gluon propagator does not change from the free propagator, as dictated by gauge invariance, so no extra dependence on $L$ is introduced by this operation.
If the cutoff length is sufficiently larger than the dynamical scale, the amplitude should be independent of the cutoff. We assume that the cutoff dependence disappears from the final result if $L\gg 1$, i.e., if the cutoff length is much larger than the string scale $1/\sqrt{\alpha}$. This means that after summing over the infinite series for each $n$, the cutoff $L$ can be effectively replaced with an order 1 number (which will be taken to be 1 below for simplicity); for $L\gg 1$, we will have
$\mathcal{F}^{(n)}(\bar{\alpha},L)\sim\mathcal{F}^{(n)}(\bar{\alpha}).$ (5.11)
Substituiting (5.10) and (5.11) into (5.9) gives,
${\cal{A}}$ $=$ $\sum_{n=0}^{\infty}\left(\frac{\bar{\alpha}}{N}\right)^{n}\mathcal{F}^{(n)}(\bar{\alpha})$ (5.12)
$=$ $\sum_{n=0}^{\infty}\bar{g}^{2n}\mathcal{F}^{(n)}(\bar{\alpha}),$ (5.14)
assuming the cutoff length is sufficiently large.
The flat-space limit is obtained by letting $\bar{\alpha}\to\infty$ with $\bar{g}$ kept finite. For this limit to make sense—we assume that it does— the functions $\mathcal{F}^{(n)}(\bar{\alpha})$ must have finite limits [15],
$\lim_{\bar{\alpha}\to\infty}\mathcal{F}^{(n)}(\bar{\alpha})={\cal{F}}_{n}.$ (5.15)
@@
${\cal{A}}=\sum_{n=0}^{\infty}{\cal{F}}_{n}\bar{g}^{2n}.$ (5.16)
Equation (5.16) has the following interpretation: The sum is over worldsheets of genus $n/2.$ The factors ${\cal{F}}_{n}$ are the worldsheet amplitudes for genus $n/2.$ And most interestingly the factor $\bar{g}^{2n}$ is the dependence on the open string coupling constant $g_{st}$ for a worldsheet of genus $n/2$, with the identification,
$g_{st}=\bar{g}.$ (5.17)
We note that the only dimensionless parameter in the flat space limit of $QCD_{2}$ is the open string coupling $g_{st}.$
@@ 5.2 DSSYK-QCD2 Correspondence
$\rm DSSYK_{\infty}$} also has a $1/N$ expansion with essentially the same form as (5.9) but with $\bar{\alpha}$ being replaced by $p^{2}$ [15]
${\cal{A}}=\sum_{n}\frac{\tilde{\mathcal{F}}^{(n)}(p^{2})}{N^{n}}.$ (5.18)
The duality between $\rm DSSYK_{\infty}$} and $QCD_{2}$ requires that we make the identification [^12] This identification of the parameters is different from the one adopted in the previous paper by two of the present authors [15], in which the similarity of perturbation expansions in SYK and QCD has been discussed. In that paper, QCD with degrees of freedom in the adjoint representation was considered, and the authors found the identification $p=\alpha$. In that case, $N$ in SYK corresponds to $N_{ym}^{2}$ in QCD, as naturally expected from the matching of the number of the degrees of freedom, which leads to the identification $\lambda=g_{ym}^{4}$.
$p^{2}=\bar{\alpha}.$ (5.19)
Dividing by $N$ and recalling that $\frac{p^{2}}{N}=\lambda,$ we find the correspondence,
$\lambda=\bar{g}^{2}=g_{st}^{2}.$ (5.20)
Thus $\lambda$ corresponds to the square of the open string coupling.
@@ 5.3 The Meson Tower
The ’t Hooft model is the large $N$ limit of a gauge theory, but it may also be thought of as an open string theory. The spectrum of mesons in the model consists of a single tower or Regge-like trajectory. Figure 7 is taken from [14] and shows the Regge trajectory [^13] See [17, 18, 19] for recent work on the mass spectrum in the ’t Hooft model. We thank Pavel Meshcheriakov for discussion on these results. for a particular non-zero value of $m_{q}.$ The horizontal axis is the meson mass squared; the vertical axis is an integer $n\geq 2$ analogous to angular momentum in higher dimensions. Asymptotically for large $n$ the trajectory approaches a straight line.
@@
The operators in (5.21) are in a sense complete in that any gauge invariant operator is a function of them. In addition to the single meson states there are multiple meson states and these span the space of states of the ’t Hooft model.
@@
We will come back to this in a future publication.
6 QCD in Rindler Space and DSSYK